OpenStudy (anonymous):

What are the solutions of 4x2 − 4x − 3 = 0 ? Need help please x = −1/2 , x = 3/2 x =1/2 , x = −3/2 x =1/4 , x = −3 x = −1/4, x = 3

OpenStudy (saifoo.khan):

\[\Huge A.\]

OpenStudy (anonymous):

Is it the first one?

OpenStudy (saifoo.khan):

Yesh!

OpenStudy (anonymous):

Okay thamkkkss :)

OpenStudy (saifoo.khan):

welks Megan Foxx.

OpenStudy (saifoo.khan):

;]

OpenStudy (anonymous):

hah :)

OpenStudy (anonymous):

x = −1/2 , x = 3/2